Please help me solve: use identities to find sin2x given cosx3/4 and tan>0
Answer by Alan3354(69443)

use identities to find sin2x given cosx3/4 and tan>0
------------
Do you mean cos(x) = 3/4 ?
If so,
sin(x) = sqrt(1 - cos^2) = sqrt(7/16) = sqrt(7)/4
----
sin(2x) = 2*sin(x)*cos(x)
= 2*(3/4)*sqrt(7)/4
= 3sqrt(7)/8